OpenStudy (anonymous):

Solve 2/(x-1) + 3/(y+5) = 11/2 -1/(x-1) + 2/(y+5)= -1

OpenStudy (anonymous):

It's a system. How do you solve it.

OpenStudy (callisto):

2/(x-1) + 3/(y+5) = 11/2 -(1) -1/(x-1) + 2/(y+5)= -1 -(2) Try (2)x2 + (1), see what you will get..

OpenStudy (anonymous):

Seiga I feel like that is wrong.

OpenStudy (anonymous):

Callisto your sentence is unclear. Is that an x^2

OpenStudy (callisto):

I meant multiply the second equation by 2, the add it to the first equation.

OpenStudy (anonymous):

No that method would yield values like xy and such and that is too complex for this remedial math class my friend is in.

OpenStudy (anonymous):

I could try bringing the denominator up.

OpenStudy (anonymous):

Or merging the two.

OpenStudy (callisto):

Let see...